OpenStudy (anonymous):

Jon earned $38,000 last year, He paid 6,840 towards entertainment. What percent of his earnings did Jon pay in entertainment expenses?

OpenStudy (anonymous):

18%

OpenStudy (anonymous):

6840=x(3800) 6840/1000=x 0.18=x Then multiply by 100 since you are for a percentage. 0.18(100)=x 18%=x

OpenStudy (anonymous):

Sorry, since you are looking for a percentage

OpenStudy (anonymous):

Remember to word these percentage problem like this : 6840 is what percent of 38000.

OpenStudy (anonymous):

Sorry again, I'm on my phone! Let me rewrite that : 6840=x(38000) 6840/38000=x 0.18=x 0.18(100)=x 18%=x

OpenStudy (anonymous):

Got it?